### v2

# Pivoting Factorization: A Compact Meta Low-Rank Representation of Sparsity for Efficient Inference in Large Language Models

Abstract
The rapid growth of Large Language Models has driven demand for effective model compression techniques to reduce memory and computation costs. Low-rank pruning has gained attention for its GPU compatibility across all densities. However, low-rank pruning struggles to match the performance of semi-structured pruning, often doubling perplexity at similar densities. In this paper, we propose Pivoting Factorization (PIFA), a novel lossless meta low-rank representation that unsupervisedly learns a compact form of any low-rank representation, effectively eliminating redundant information. PIFA identifies pivot rows (linearly independent rows) and expresses non-pivot rows as linear combinations, achieving 24.2% additional memory savings and 24.6% faster inference over low-rank layers at rank = 50% of dimension. To mitigate the performance degradation caused by low-rank pruning, we introduce a novel, retraining-free reconstruction method that minimizes error accumulation (M). MPIFA, combining M and PIFA into an end-to-end framework, significantly outperforms existing low-rank pruning methods, and achieves performance comparable to semi-structured pruning, while surpassing it in GPU efficiency and compatibility. Our code is available at https://github.com/biomedical-cybernetics/pivoting-factorization.
[…]
tableComparison of PIFA with other sparsity. Method CPU Speedup GPU Speedup GPU Mem Reduction Any Sparsity GPU Support Performance Unstructured Sparsity ✓✓✓ Semi-Structured Sparsity Ampere GPU ✓✓ Structured Sparsity General SVD-Based Low-Rank Sparsity General PIFA Low-Rank Sparsity General ✓✓
Figure 1: Comparison of parameter ratios.
1 Introduction
Figure 2: Illustration of the low-rank pruning method MPIFA (Algorithm 3), which consists of: (a) Online Error-Accumulation-Minimization Reconstruction (M). Block $R$ solves the least-squares optimization problem. The improvements upon SVD-LLM’s full-batch reconstruction, highlighted in red, include using the dense data flow to minimize error accumulation, and processing each sample sequentially to avoid GPU memory overflow. (b) Pivoting Factorization (PIFA). For any singular matrix with rank $r$, Pivoting Factorization further reduces $r^{2}-r$ parameters, with no additional loss induced.
The rapid growth of Large Language Models (LLMs) (Radford, 2018; Radford et al., 2019; Mann et al., 2020; Touvron et al., 2023a) has revolutionized natural language processing tasks but has also introduced significant challenges related to memory consumption and computational costs. Deploying these models efficiently, particularly on resource-limited hardware, has driven a surge of interest in model compression techniques (Wan et al., 2023; Zhu et al., 2024). Among these techniques, semi-structured pruning, specifically N:M sparsity, has emerged as a promising approach due to its hardware-friendly nature, enabling efficient acceleration on NVIDIA’s Ampere GPUs (Mishra et al., 2021; nvi, 2020). However, semi-structured pruning suffers from two major limitations: it is restricted to specific hardware architectures, and it couldn’t flexible adjust density.
In contrast, low-rank pruning methods, primarily based on Singular Value Decomposition (SVD), preserve the coherence of tensor shapes, making them universally compatible with any GPU architecture at any density. Recent works (Yuan et al., 2023; Wang et al., 2024) have demonstrated the potential of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods struggle to compete with semi-structured pruning in terms of performance, often resulting in a 2x increase in perplexity (PPL) at the same densities. This performance gap stems primarily from two challenges: (1) Low-rank pruning introduces information redundancy in the decomposed weight matrices, and (2) existing reconstruction methods accumulate errors across layers, leading to suboptimal performance.
To address challenge (1), we propose Pivoting Factorization (PIFA), a novel lossless low-rank representation that eliminates redundancy and enhances computational efficiency. PIFA is a meta low-rank representation, because it unsupervisedly learns a compact representation of any other learned low-rank representation. PIFA identifies $r$ linearly independent rows from a singular weight matrix, which we refer to as pivot rows, and represents all other rows as linear combinations of these pivot rows. PIFA achieves significant improvements in both speedup and memory reduction during inference. Specifically, at $r/d=0.5$, the PIFA layer achieves 24.2% additional memory savings and 24.6% faster inference compared to SVD-based low-rank layers, without inducing any loss.
To address challenge (2), we propose an Online Error-Accumulation-Minimization Reconstruction (M) algorithm that minimizes error accumulation—a problem pervasive in existing reconstruction methods for both low-rank pruning (Wang et al., 2024) and semi-structured pruning (Frantar & Alistarh, 2023; Li et al., 2024). Existing methods rely on degraded data flow, where accumulated errors from previous modules propagate through the reconstruction process, leading to suboptimal performance. Our approach addresses this issue by combining dense data flow and low-rank data flow, as reconstruction targets, effectively mitigating the errors carried forward from earlier layers. Furthermore, the method operates online, processing large numbers of calibration samples sequentially to stay within GPU memory constraints.
Combining M and PIFA, we present MPIFA—an end-to-end, retraining-free low-rank pruning framework for LLMs. MPIFA significantly outperforms existing low-rank pruning methods, reducing the perplexity gap by 40%-70% on LLaMA2 models (7B, 13B, 70B) and the LLaMA3-8B model  [^1] https://www.llama.com/. Further experiments show that MPIFA consistently achieves superior speedup and memory savings both layerwise and end-to-end compared to semi-structured pruning, while maintaining comparable or even better perplexity. Notably, as shown in Table 3, at $d=32768$, PIFA with 55% density achieves a 2.1$\times$ speedup, whereas various implementations of semi-sparse methods are either slower than the dense linear layer or fail to execute.
The two main contributions of this work can be summarized as follows:
  1. 1.
    We propose Pivoting Factorization (PIFA), a novel lossless meta low-rank representation that unsupervisedly learns a compact form of any SVD-based low-rank representation, effectively compressing out redundant information.
  2. 2.
    We introduce an Online Error-Accumulation-Minimization Reconstruction (M) algorithm that mitigates error accumulation by leveraging multiple data flows for reconstruction.
2 Related Work
2.1 Connection-wise pruning
Pruning methods
We define connection-wise pruning as removing certain connections between neurons in the network that are deemed less important. To achieve this, a series of methods have been proposed. Optimal Brain Damage (OBD) (Le Cun et al., 1990) and Optimal Brain Surgeon (OBS) (Hassibi et al., 1993) were proposed to identify the weight saliency by computing the Hessian matrix using calibration data. Recent methods such as SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), RIA (Zhang et al., 2024), along with other works (Fang et al., 2024; Dong et al., 2024; Das et al., 2024), have advanced these ideas. Wanda prunes weights with the smallest magnitudes multiplied by input activations. Relative Importance and Activations (RIA) jointly considers both the input and output channels of weights along with activation information. Furthermore, OWL (Yin et al., ) explores non-uniform sparsity by pruning based on the distribution of outlier activations, while other works (Lu et al., 2024; Mocanu et al., 2018; Ye et al., 2020; Zhuang et al., 2018) investigate alternative criteria for non-uniform sparsity.
Pruning granularity (compared in Table 1):
  1. 1.
    Unstructured pruning removes individual weights based on specific criteria. Today, unstructured pruning is a critical technique for compressing large language models (LLMs) to balance performance and computational efficiency. However, unstructured pruning can only accelerate computations on CPUs due to its unstructured sparsity pattern.
  2. 2.
    Semi-structured pruning, i.e., N:M sparsity, enforces that in every group of $M$ consecutive elements, $N$ must be zeroed out. This constraint is hardware-friendly and enables optimized acceleration on GPUs like NVIDIA’s Ampere architecture (Mishra et al., 2021). However, semi-structured pruning is constrained by its sparsity pattern, preventing flexible density adjustments and making it inapplicable for acceleration on general GPUs.
  3. 3.
    Structured pruning (Ma et al., 2023; van der Ouderaa et al., 2024; Ashkboos et al., 2024; Lin et al., 2024; Song et al., 2024; Men et al., 2024; Gao et al., 2024) removes entire components of the model, such as neurons, channels, or attention heads, rather than individual weights. This method preserves tensor alignment and coherence, ensuring compatibility with all GPUs and enabling significant acceleration on both CPUs and GPUs. However, in LLMs, structured pruning can lead to greater loss compared to unstructured or semi-structured pruning.
2.2 Low-rank pruning
Low-rank pruning applies matrix decomposition techniques, such as Singular Value Decomposition (SVD), to approximate weight matrices with lower-rank representations, thereby reducing both storage and computational demands. This approach, compatible with any GPU, represents large matrices as products of smaller ones, improving computational efficiency. Recent studies (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024; Jaiswal et al., 2024; Saha et al., 2024; Kaushal et al., 2023; Sharma et al., 2023; Qinsi et al., ) highlight the effectiveness of low-rank decomposition in compressing LLMs. However, despite their flexibility, these methods lag behind semi-structured pruning in performance, often leading to a 2$\times$ increase in perplexity (PPL) at the same densities.
3 Lossless Low-Rank Compression
3.1 Motivation: Information Redundancy in Singular Value Decomposition
For a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, Low-rank pruning methods (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024) decompose the matrix into a product of two low-rank matrices, $\mathbf{W}\approx\vec{U}\vec{V}^{\mathrm{T}}$, where $\vec{U}\in\mathbb{R}^{m\times r}$ and $\vec{V}^{\mathrm{T}}\in\mathbb{R}^{r\times n}$, forming a low-rank approximation of $\mathbf{W}$. Consider naive SVD pruning as an example. First, the weight matrix is factorized using SVD as $\mathbf{W}=\vec{B}\vec{E}\vec{A}^{\mathrm{T}}$. Next, the top-$r$ singular values and corresponding singular vectors are retained, denoted as $\vec{B}_{r}$, $\vec{E}_{r}$, and $\vec{A}_{r}^{\mathrm{T}}$. Finally, the singular values $\vec{E}_{r}$ are merged into the singular vectors, yielding $\vec{U}=\vec{B}_{r}\vec{E}_{r}$ and $\vec{V}^{\mathrm{T}}=\vec{A}_{r}^{\mathrm{T}}$.
The number of parameters in the dense weight matrix is $mn$, whereas the total number of parameters in the low-rank matrices is $r(m+n)$. As shown in Figure 1, SVD-based low-rank representations fail to compress the weight matrix when $r$ exceeds half of the matrix dimensions. However, in low-rank decomposition, the orthogonality constraints among singular vectors reduce the effective degrees of freedom. Specifically, for $\vec{U}$ and $\vec{V}^{\mathrm{T}}$, there are $\binom{r}{2}=\frac{r(r-1)}{2}$ unique pairs of singular vectors for each matrix, and each pair imposes one linear constraint due to orthogonality. Together, these constraints reduce the total degrees of freedom by $r(r-1)$.
Thus, the actual degrees of freedom in the low-rank representation are $r(m+n)-(r^{2}-r)$. This reveals redundancy in low-rank representations and suggests the possibility of encoding the same information with fewer parameters by eliminating it.
Question: Can we design a matrix factorization method that reduces parameters to $r(m+n)-(r^{2}-r)$ without losing representational power?
3.2 Pivoting Factorization
To address the previously discussed question, we propose Pivoting Factorization (PIFA), a novel matrix factorization method. For any low-rank matrices, which can be obtained by any low-rank pruning methods, PIFA further reduces parameters without inducing additional loss.
The process of Pivoting Factorization is illustrated in Figure 2(b). Given a weight matrix already decomposed into low-rank matrices $\vec{U}\in\mathbb{R}^{m\times r}$ and $\vec{V}^{\mathrm{T}}\in\mathbb{R}^{r\times n}$, we first multiply $\vec{U}$ and $\vec{V}^{\mathrm{T}}$ to get the singular matrix, $\mathbf{W}^{\prime}=\vec{U}\vec{V}^{\mathrm{T}}$. Since $\mathbf{W}^{\prime}$ has rank $r$, it contains $r$ linearly independent rows, also referred to as pivot rows. The set of linearly independent rows can be identified using LU or QR decomposition with pivoting (Businger & Golub, 1971). Let $\mathcal{I}$ represent the set of row indices corresponding to the $r$ pivot rows of $\mathbf{W}^{\prime}$. Thus, any non-pivot row can be expressed as a linear combination of these $r$ pivot rows. Denoting the set of non-pivot row indices as $\mathcal{I}^{c}=\{1,2,...,m\}\backslash\mathcal{I}$, then we define:
$\mathbf{W}_{p}=\mathbf{W}^{\prime}[\mathcal{I},:],\quad\mathbf{W}_{np}=\mathbf{W}^{\prime}[\mathcal{I}^{c},:]$ (1)
where $\mathbf{W}_{p}\in\mathbb{R}^{r\times n}$ is the pivot-row matrix, and $\mathbf{W}_{np}\in\mathbb{R}^{{(m-r)}\times n}$ is the non-pivot-row matrix. With the definition of non-pivot rows, non-pivot-row matrix can be expressed as:
$\mathbf{W}_{np}=\mathbf{C}\mathbf{W}_{p}$ (2)
where $\mathbf{C}\in\mathbb{R}^{{(m-r)}\times r}$ is the coefficient matrix. Algorithm 1 details the PIFA process, which generates the components of a PIFA layer: pivot-row indices $\mathcal{I}$, the pivot-row matrix $\mathbf{W}_{p}$, and the coefficient matrix $\mathbf{C}$. Algorithm 2 describes the inference procedure for the PIFA layer, which leverages $\mathbf{W}_{p}$, $\mathbf{C}$ and $\mathcal{I}$ to compute the output.
Algorithm 1 Pivoting Factorization
0:  Low-rank matrix $\mathbf{W}^{\prime}\in\mathbb{R}^{m\times n}$ with rank $r$
1:  Use QR (or LU) decomposition with pivoting to find pivot-row indices: $\mathcal{I}\leftarrow\mathrm{QR}_{\text{pivot}}(\mathbf{W}^{\prime})$
2:  Define $\mathcal{I}^{c}=\{1,2,\dots,m\}\setminus\mathcal{I}$, the complement of $\mathcal{I}$, representing non-pivot row indices
3:  Compute pivot-row matrix: $\mathbf{W}_{p}\leftarrow\mathbf{W}^{\prime}[\mathcal{I},:]$
4:  Compute non-pivot-row matrix: $\mathbf{W}_{np}\leftarrow\mathbf{W}^{\prime}[\mathcal{I}^{c},:]$
5:  Compute coefficient matrix $\mathbf{C}$ by solving matrix equation: $\mathbf{C}\leftarrow\mathrm{linsolve}(\mathbf{W}_{np}=\mathbf{C}\mathbf{W}_{p})$
5:  PIFA layer $P$: 1) Pivot-row indices $\mathcal{I}\in\mathbb{R}^{r}$; 2) pivot-row matrix $\mathbf{W}_{p}\in\mathbb{R}^{r\times n}$; 3) coefficient matrix $\mathbf{C}\in\mathbb{R}^{(m-r)\times r}$ for non-pivot rows
Algorithm 2 PIFA Layer
0:  Input $\mathbf{X}\in\mathbb{R}^{n\times b}$, where $b$ is batch size; pivot-row indices $\mathcal{I}\in\mathbb{R}^{r}$; pivot-row matrix $\mathbf{W}_{p}\in\mathbb{R}^{r\times n}$; coefficient matrix $\mathbf{C}\in\mathbb{R}^{(m-r)\times r}$ for non-pivot rows
1:  Define $\mathcal{I}^{c}=\{1,2,\dots,m\}\setminus\mathcal{I}$ representing non-pivot row indices
2:  Compute output of pivot channels: $\mathbf{Y}_{p}\leftarrow\mathbf{W}_{p}\mathbf{X}$
3:  Compute output of non-pivot channels: $\mathbf{Y}_{np}\leftarrow\mathbf{C}\mathbf{Y}_{p}$
4:  Assign pivot channels to output: $\mathbf{Y}[\mathcal{I},:]\leftarrow\mathbf{Y}_{p}$
5:  Assign non-pivot channels to output: $\mathbf{Y}[\mathcal{I}^{c},:]\leftarrow\mathbf{Y}_{np}$
5:  $\mathbf{Y}$
3.3 Memory and Computational Cost of PIFA
Memory cost of PIFA.
For each low-rank weight matrix $\mathbf{W^{\prime}}$, PIFA needs to store $\mathcal{I}$, $\mathbf{W}_{p}$ and $\mathbf{C}$, totaling $r(m+n)-r^{2}+r$. Figure 1 illustrates the relationship between the number of parameters in PIFA, traditional low-rank decomposition, and a dense weight matrix (square).
Since $r(m+n)>r(m+n)-r^{2}+r$ for any rank $r$, PIFA consistently requires less memory than traditional low-rank decomposition. For the comparison with dense weight matrix, because $r<\min(m,n)$, we have:
$(m-r)(n-r)>0\quad\Rightarrow\quad mn>r(m+n)-r^{2}$ (3)
Neglecting the pivot-row index $\mathcal{I}$, which has negligible memory overhead compared to other variables, PIFA could consistently consumes less memory than a dense weight matrix. In contrast, traditional low-rank decomposition may exceed the memory cost of dense matrices when $r>\frac{mn}{m+n}$.
Computational cost of PIFA.
We analyze the computational cost of each linear layer for an input batch size $b$, where $\mathbf{X}\in\mathbb{R}^{n\times b}$. We compute the FLOPs for each method as follows:
  • For the dense linear layer $\mathbf{Y}=\mathbf{W}\mathbf{X}$, where $\mathbf{W}\in\mathbb{R}^{m\times n}$ and $\mathbf{X}\in\mathbb{R}^{n\times b}$, the computational cost is $2mnb$ FLOPs.
  • For the traditional low-rank layer $\mathbf{Y}=\vec{U}\vec{V}^{\mathrm{T}}\mathbf{X}$, where $\vec{U}\in\mathbb{R}^{m\times r}$ and $\vec{V}^{\mathrm{T}}\in\mathbb{R}^{r\times n}$, the computational cost includes: $\vec{V}^{\mathrm{T}}\mathbf{X}$ ($2rnb$) and $\vec{U}(\vec{V}^{\mathrm{T}}\mathbf{X})$ ($2mrb$). The total cost is $2rnb+2mrb=2br(m+n)$ FLOPs.
  • For the PIFA layer (Algorithm 2), the computational cost includes: $\mathbf{Y}_{p}\leftarrow\mathbf{W}_{p}\mathbf{X}$ $(2rnb)$ and $\mathbf{Y}_{np}\leftarrow\mathbf{C}\mathbf{Y}_{p}$ $2br(m-r)$. The total cost is $2rnb+2br(m-r)=2br(m+n-r)$ FLOPs.
PIFA’s computational cost is proportional to its memory cost, differing only by a factor of $2b$. As a result, PIFA consistently outperforms both dense linear layers and traditional low-rank layers in computational efficiency.
Figure 3: Pivoting Factorization vs. LU and QR decompositions. Applied to the permuted matrix (pivot rows at the top), Pivoting Factorization avoids the trapezoidal distribution of non-trivial parameters in LU decomposition, instead reorganizing them into a rectangular pattern. This structure optimizes GPU memory usage and reduces computation overhead.
Comparison with LU and QR decomposition.
Figure 3 compares the structure of LU and QR decomposition with Pivoting Factorization on a permuted weight matrix, where pivot rows have already been moved to the top. LU decomposition retains the same number of non-trivial parameters (i.e., those not preset as zero or one) as Pivoting Factorization. However, the trapezoidal distribution of non-trivial parameters in LU decomposition complicates efficient storage and computation on the GPU. In contrast, PIFA reorganizes all non-trivial parameters into a rectangular distribution, which is more GPU-friendly for storage and computation. Thus, Pivoting Factorization is more efficient for GPU computation.
4 Online Error-Accumulation-Minimization Reconstruction
In addition to PIFA, we propose a novel Online Error-Accumulation-Minimization Reconstruction (M) method (illustrated in Figure 2(a)) to reconstruct the low-rank matrix before applying PIFA.
A low-rank pruning step is first required to obtain the low-rank matrices $\vec{U}$ and $\vec{V}^{\mathrm{T}}$ before reconstruction. To achieve this, we adopt the pruning method from SVD-LLM (Wang et al., 2024), which has demonstrated superior performance among existing methods.
SVD-LLM first introduced low-rank matrix reconstruction. It updates $\mathbf{U}$ using a closed-form least squares solution:
$\mathbf{U}_{r}$ $=\arg\min_{\mathbf{U}}\|\mathbf{W}\mathbf{X}-\vec{U}\vec{V}^{\mathrm{T}}\mathbf{X}\|_{\text{F}}$ (4)
$=\mathbf{W}\mathbf{X}\mathbf{D}^{\mathrm{T}}(\mathbf{D}\mathbf{D}^{\mathrm{T}})^{-1},\mathbf{D}=\vec{V}^{\mathrm{T}}\mathbf{X}$
where $\mathbf{X}$ is the calibration data. We improve Equation 4 in the following aspects:
① Online algorithm.
Equation 4 requires loading the entire calibration dataset $\mathbf{X}$ into GPU memory to compute the least squares solution. As a result, the number of calibration samples is limited to a maximum of 16 on LLaMA2-7B (4 on LLaMA2-70B) with a 48GB A6000 GPU, leading to overfitting to the calibration data (see Section 5.3).
Applying the associative property of matrix multiplication, we reformulate Equation 4 into its online version:
$\mathbf{U}_{r}=\mathbf{W}(\mathbf{X}\mathbf{X}^{\mathrm{T}})\vec{V}(\vec{V}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})\vec{V})^{-1}$ (5)
The term $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ can be computed incrementally as $\mathbf{X}\mathbf{X}^{\mathrm{T}}=\sum_{i=1}^{b}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathrm{T}}$, where $\mathbf{x}_{i}$ represents the input of sample $i$. As $\mathbf{X}\mathbf{X}^{\mathrm{T}}\in\mathbb{R}^{n\times n}$, the memory consumption of the online least squares solution remains constant, regardless of the number of calibration samples.
② Error Accumulation Minimization.
Existing reconstruction methods in low-rank pruning (Wang et al., 2024) and semi-structured pruning (Frantar & Alistarh, 2023; Li et al., 2024) rely solely on one data flow, i.e. low-rank data flow in Figure 2. This approach allows accumulated errors from previous modules to propagate through the reconstruction process, potentially degrading performance, as each subsequent module is optimized based on an already-degraded data flow.
Our method mitigates this issue by correcting accumulated error at each module, realigning it with the dense data flow:
$\min\|\mathbf{W}\mathbf{X}_{o}-\vec{U}{\vec{V}}^{\mathrm{T}}\mathbf{X}_{u}\|_{\text{F}}$ (6)
where $\mathbf{X}_{o}$ represents the dense data flow input, produced by the previous layer’s dense weight, and $\mathbf{X}_{u}$ represents the low-rank data flow input, produced by the previous layer’s low-rank weight. This ensures that each module’s output remains aligned with the output of original model, recovering the accumulated error in $\mathbf{X}_{u}$.
However, in practical experiments, we observe that relying solely on the dense data flow output $\mathbf{W}\mathbf{X}_{o}$ as the reconstruction target tends to overfit the reconstructed low-rank weight to the distribution of the calibration data. Using a combination of dense and low rank data flow outputs mitigates overfitting:
$\mathbf{Y}_{t}=\lambda\mathbf{W}\mathbf{X}_{o}+(1-\lambda)\mathbf{W}\mathbf{X}_{u}$ (7)
where $\lambda$ is the mix ratio. The optimization target becomes $\min\|\mathbf{Y}_{t}-\vec{U}{\vec{V}}^{\mathrm{T}}\mathbf{X}_{u}\|_{\text{F}}$. The low-rank data flow output serves as a regularization term, minimizing the distance between $\vec{U}{\vec{V}}^{\mathrm{T}}\mathbf{X}_{u}$ and $\mathbf{W}\mathbf{X}_{u}$. Since $\mathbf{W}$ has been optimized on a much larger and more diverse pre-training dataset, using it as a constraint helps $\vec{U}{\vec{V}}^{\mathrm{T}}$ generalize better and prevents overfitting to the limited calibration data. Empirically, we found that the mix ratio $\lambda=0.25$ achieves the best performance (see ablation study in Section 5.3).
③ Reconstructing both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$.
Equation 4 reconstructs only $\mathbf{U}$. We find it beneficial to also update $\vec{V}^{\mathrm{T}}$and provide the closed-form solution:
$\mathbf{V}_{r}^{\mathrm{T}}$ $=\arg\min_{\mathbf{V}^{\mathrm{T}}}\|\mathbf{Y}_{t}-\vec{U}\vec{V}^{\mathrm{T}}\mathbf{X}\|_{\text{F}}$ (8)
$=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{Y}_{t}\mathbf{X}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}$
The proof is provided in Appendix A. Updating $\vec{V}^{\mathrm{T}}$ can also be performed online by incrementally computing $\mathbf{Y}\mathbf{X}^{\mathrm{T}}$ and $\mathbf{X}\mathbf{X}^{\mathrm{T}}$.
[…]
tablePerplexity ($\downarrow$) at different parameter density (proportion of remaining parameters relative to the original model) on WikiText2. The best-performing method is highlighted in bold. Density Model Method 100% 90% 80% 70% 60% 50% 40% LLaMA2-7B SVD 5.47 16063 18236 30588 39632 53179 65072 ASVD 5.91 9.53 221.6 5401 26040 24178 SVD-LLM 7.27 8.38 10.66 16.11 27.19 54.20 MPIFA 5.69 6.16 7.05 8.81 12.77 21.25 LLaMA2-13B SVD 4.88 2168 6177 37827 24149 14349 41758 ASVD 5.12 6.67 17.03 587.1 3103 4197 SVD-LLM 5.94 6.66 8.00 10.79 18.38 42.79 MPIFA 5.03 5.39 7.12 7.41 10.30 16.72 LLaMA2-70B SVD 3.32 6.77 17.70 203.7 2218 6803 15856 ASVD OOM OOM OOM OOM OOM OOM SVD-LLM 4.12 4.58 5.31 6.60 9.09 14.82 MPIFA 3.54 3.96 4.58 5.54 7.40 12.01 LLaMA3-8B SVD 6.14 463461 626967 154679 62640 144064 216552 ASVD 9.37 275.6 12553 21756 185265 13504 SVD-LLM 9.83 13.62 23.66 42.60 83.46 163.5 MPIFA 6.93 8.31 10.83 16.41 28.90 47.02
5 Experiments
MPIFA.
We combine Online Error-Accumulation-Minimization Reconstruction with Pivoting Factorization into an end-to-end low-rank compression method, MPIFA (illustrated in Figure 2). MPIFA proceeds as follows: ① First, Online Error-Accumulation-Minimization Reconstruction is applied to obtain and refine the low-rank matrices $\vec{U}_{r}$ and $\vec{V}_{r}^{\mathrm{T}}$; ② Then, PIFA decomposes the singular matrix $\mathbf{W}^{\prime}=\vec{U}_{r}\vec{V}_{r}^{\mathrm{T}}$ into $\mathcal{I},\mathbf{W}_{p},\mathbf{C}\leftarrow\mathrm{PIFA}(\mathbf{W}^{\prime})$, which are stored in a PIFA layer that replaces the original linear layer.
MPIFANS.
Compared to semi-structured sparsity, MPIFA offers the advantage of allowing non-uniform sparsity for each module. We term the Non-uniform Sparsity version of MPIFA as MPIFANS. The detailed implementation of MPIFANS can be found in Appendix B.2.
Models and Datasets.
We apply MPIFA to pre-trained LLMs: LLaMA2 (7B, 13B, 70B) (Touvron et al., 2023b) and LLaMA3 (8B) (Dubey et al., 2024). Both the calibration data and perplexity (PPL) evaluations are based on the WikiText2 dataset (Merity et al., 2022), with a sequence length of 2048 tokens for all experiments.
5.1 Main Result
Comparison with other low-rank pruning.
We evaluate MPIFA against state-of-the-art low-rank pruning methods: ASVD (Yuan et al., 2023) and SVD-LLM (Wang et al., 2024). Vanilla SVD is also included for reference. SVD-LLM has two versions, as detailed in their original paper. Following their approach, we select the best-performing version for each density. Results for both versions of SVD-LLM are provided in Table 5.3. MPIFA utilizes 128 calibration samples and sets $\lambda=0.25$. For all models except LLaMA-2-70B, MPIFA reconstructs both $\vec{U}$ and $\vec{V}^{\mathrm{T}}$. For LLaMA-2-70B, MPIFA reconstructs only $\vec{U}$.
Table 4 displays the test perplexity (PPL) of each low-rank pruning method on WikiText2 under 0.4-0.9 density, which is defined as the proportion of parameters remaining compared with the original model. The results show that MPIFA significantly outperforms other low-rank pruning method, reducing the perplexity gap by 66.4% (LLaMA2-7B), 53.8% (LLaMA2-13B), 40.7% (LLaMA2-70B), and 72.7% (LLaMA3-8B) on average. The perplexity gap is defined as the difference between the perplexity of a low-rank pruning method and the original model.
Comparison with semi-structured pruning.
We further compare MPIFA with 2:4 semi-structured pruning methods: magnitude pruning (Zhu & Gupta, 2017), and two recent state-of-the-art works Wanda (Sun et al., 2024), and RIA (Zhang et al., 2024). According to (Mishra et al., 2021), for 16-bit operands, 2:4 sparse leads to $\sim$44savings in GPU memory. Therefore, we compare 2:4 sparse method with MPIFA at 0.55 density to ensure that all methods achieve the same memory reduction (see Table 3 for memory comparison).
Table 1 shows that MPIFANS outperforms 2:4 pruning methods, reducing the perplexity gap by 21.7% for LLaMA2-7B and 3.2% for LLaMA2-13B.
Table 1: Perplexity ($\downarrow$) comparison with semi-structured pruning under the same memory reduction on WikiText2. The best performance pruning method is indicated in bold. MPIFANS means MPIFA using non-uniform sparsity.
Method LLaMA2-7B LLaMA2-13B
Dense 5.47 4.88
Magnitude 2:4 37.77 8.89
Wanda 2:4 11.40 8.33
RIA 2:4 10.85 8.03
SVD 55% 69128 24947
ASVD 55% 9370 2039
SVD-LLM 55% 20.43 13.69
MPIFANS 55% 9.68 7.93
Fine-tuning After Pruning.
We investigate how fine-tuning helps recover the performance loss caused by low-rank pruning. The detailed experimental settings are provided in Appendix B.3. As shown in Table 2, fine-tuned MPIFANS achieves the best performance among all fine-tuned pruning methods, bringing its perplexity close to the dense baseline at 55% density. Unlike semi-structured methods, which cannot accelerate the backward pass due to transposed weight tensors violating the 2:4 constraint (Mishra et al., 2021), PIFA and other low-rank methods enable acceleration in both the forward and backward passes.
Table 2: Perplexity ($\downarrow$) of pruned models after fine-tuning on WikiText2 (LLaMA2-7B). The best-performance pruning method is indicated in bold.
Method LLaMA2-7B
Dense 5.47
Magnitude 2:4 6.63
Wanda 2:4 6.40
RIA 2:4 6.37
SVD 55% 9.24
ASVD 55% 8.64
SVD-LLM 55% 7.36
MPIFANS 55% 6.34
Figure 4: Layerwise speedup of the PIFA layer and semi-sparse layers across various dimensions, compared to dense linear layers on the same GPU, with a sequence length of 2048 and a batch size of 32, using FP16 (FP32 is not supported by 2:4 sparsity in torch.sparse). PIFA shows increasing speedup as the dimension grows. Detailed values are provided in Table 3.
5.2 Inference Speedup and Memory Reduction
PIFA layer vs low-rank layer.
The PIFA layer achieves significant savings in both memory and computation time, as shown in Figure 7, which compares its actual runtime and memory usage with those of a standard linear layer and an SVD-based low-rank layer. In FP32, at $50\%$ density, PIFA achieves $47.6\%$ memory savings and 1.95x speedup, closely matching the ideal memory and speedup. Additionally, at the same rank, PIFA consistently achieves higher compression and faster inference than the low-rank layer. For example, at $r/d=0.5$, PIFA losslessly compresses the memory of the low-rank layer by $24.2\%$ and reduces inference time by $24.6\%$.
PIFA layer vs semi-sparse layer.
Figure 4 and Table 3 compare the speedup and memory usage of the PIFA layer with 2:4 semi-sparse linear layers. The semi-sparse layers are implemented using cuSPARSELt [^2] https://docs.nvidia.com/cuda/cusparselt/ or CUTLASS [^3] https://github.com/NVIDIA/cutlass library, with the reported speedup representing the higher value between the two implementations. The results span various dimensions on A6000 and A100 GPUs. PIFA demonstrates consistently superior efficiency, achieving the highest speedup and lowest memory usage in all configurations except $d=4096$. Notably, PIFA’s acceleration increases with matrix dimensions, reflecting its scalability and computational effectiveness. As shown in Table 3, at $d=32768$, PIFA with 55% density achieves a 2.1$\times$ speedup, while 2:4 (CUTLASS) is slower than the dense linear layer, and 2:4 (cuSPARSELt) raises an error.
End-to-end LLM inference.
Table 4 compares the end-to-end inference throughput and memory usage of MPIFANS with semi-sparsity (2:4 cuSPARSELt and CUTLASS) on LLaMA2-7B and LLaMA2-13B models in FP16. MPIFANS consistently outperforms semi-sparsity in both throughput and memory efficiency at $55\%$ density. Furthermore, the operations supported by semi-sparsity are limited in torch.sparse, resulting in errors when the KV cache is enabled, which further limits the application of semi-sparse for LLM inference.
5.3 Ablation Study
[…]
tableAblation: Impact of PIFA and M on perplexity ($\downarrow$) across parameter densities. on WikiText2. W denotes using SVD-LLM’s pruning only; W + U denotes using SVD-LLM’s pruning and full-batch reconstruction; W + M denotes using our Online Error-Accumulation-Minimization Reconstruction, which incorporates SVD-LLM’s pruning as the initial step; W + M + PIFA denotes using Online Error-Accumulation-Minimization Reconstruction followed by PIFA, i.e., MPIFA. Density Model Method 100% 90% 80% 70% 60% 50% 40% LLaMA2-7B W 5.47 7.27 8.38 10.66 16.14 33.27 89.98 W + U 7.60 8.84 11.15 16.11 27.19 54.20 W + M 6.71 7.50 8.86 11.45 16.55 25.26 W + M + PIFA (MPIFA) 5.69 6.16 7.05 8.81 12.77 21.25 LLaMA2-13B W 4.88 5.94 6.66 8.00 10.79 18.38 43.92 W + U 6.45 7.37 9.07 12.52 20.95 42.79 W + M 5.80 6.41 7.42 9.31 13.09 19.93 W + M + PIFA (MPIFA) 5.03 5.39 7.12 7.41 10.30 16.72 LLaMA2-70B W 3.32 4.12 4.58 5.31 6.60 9.09 14.82 W + U 8.23 8.33 8.66 10.02 13.41 22.39 W + M 4.15 4.63 5.31 6.46 8.72 14.11 W + M + PIFA (MPIFA) 3.54 3.96 4.58 5.54 7.40 12.01 LLaMA3-8B W 6.14 9.83 13.62 25.43 76.86 290.3 676.7 W + U 10.63 14.66 23.66 42.60 83.46 163.5 W + M 9.16 11.25 15.27 23.55 36.14 53.85 W + M + PIFA (MPIFA) 6.93 8.31 10.83 16.41 28.90 47.02
Impact of PIFA and M
Table 5.3 presents an ablation study that evaluates the impact of our Online Error-Accumulation-Minimization Reconstruction (denoted as M) and Pivoting Factorization (PIFA) on perplexity across varying parameter densities. The methods compared include:
  • W: Using only the pruning step of SVD-LLM (truncation-aware data whitening).
  • W + U: Applying SVD-LLM’s pruning followed by full-batch reconstruction.
  • W + M: Employing our Online Error-Accumulation-Minimization Reconstruction, which incorporates SVD-LLM’s pruning as the initial step.
  • W + M + PIFA: Combining Online Error-Accumulation-Minimization Reconstruction with PIFA (denoted as MPIFA).
The results reveal several key findings:
  1. 1.
    Full-batch reconstruction (W + U) occasionally worsens perplexity compared to using only the pruning step (W). This highlights the drawbacks of full-batch methods, as overfitting to the limited calibration data can degrade performance.
  2. 2.
    Our reconstruction method (W + M) consistently outperforms full-batch reconstruction (W + U) and pruning alone (W) across all models and densities. This demonstrates the effectiveness of Online Error-Accumulation-Minimization Reconstruction in reducing error accumulation and improving the compression of low-rank matrices.
  3. 3.
    PIFA further improves performance when combined with M. The W + M + PIFA configuration achieves the best perplexity across all settings, validating the advantage of applying PIFA for additional parameter reduction without inducing any additional loss.
These findings emphasize the significance of M and PIFA in achieving superior low-rank pruning performance.
Figure 5: Impact of mix ratio. With 0.5 density, MPIFA achieves lowest PPL when the mix ratio $\lambda$ in Equation 7 is around 0.25.
Impact of mix ratio $\lambda$ in M.
The mix ratio $\lambda$ in Equation 7 determines the proportion of the dense data flow in the reconstruction target. As shown in Figure 5, using a moderate ratio $\lambda=0.25$, MPIFA achieves significantly lower PPL compared to $\lambda=0$, where the reconstruction target relies solely on the low-rank data flow, as in previous studies (Wang et al., 2024; Frantar & Alistarh, 2023) did. This demonstrates the effectiveness of our error-accumulation-corrected strategy, in which the dense data flow output is beneficial as part of the reconstruction target. In Figure 5, we also observe that an excessively large $\lambda$ increases PPL, indicating overfitting to the calibration data.
(a) LLaMA2-7B
(b) LLaMA2-13B
Figure 6: Impact of calibration sample size. On MPIFA with 0.5 density, reconstructing both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ is more sensitive to the number of calibration samples than reconstructing only $\mathbf{U}$.
Impact of Calibration Sample Size.
M depends on calibration data to accurately estimate $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$. As shown in Figure 6, the perplexity of MPIFA decreases as the number of calibration samples increases. We hypothesize that increasing the number of calibration samples reduces the condition number of the least squares solution, improving numerical stability.
To investigate this, we calculate the condition numbers of $\vec{V}^{\mathrm{T}}\mathbf{X}\mathbf{X}^{\mathrm{T}}\vec{V}$ in Equation 5 and $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ in Equation 8, as their inverses are required for reconstructing $\vec{U}$ and $\vec{V}^{\mathrm{T}}$. Figure 8 presents the condition numbers for these matrices in the first layer of LLaMA2-7B. The observed reduction in condition number indicates that the matrices become less singular as the calibration sample size increases, thereby improving numerical stability when solving the least squares equations. This increased stability ultimately results in lower perplexity in the reconstructed model.
Impact of reconstructing $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$.
Figures 6a and 6b compare the effects of reconstructing only $\mathbf{U}$, only $\vec{V}^{\mathrm{T}}$, and both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ across different calibration sizes. The results indicate that with sufficient calibration samples, reconstructing both $\mathbf{U}$ and $\vec{V}^{\mathrm{T}}$ achieves lower perplexity than reconstructing only $\mathbf{U}$ or only $\vec{V}^{\mathrm{T}}$.
6 Conclusion and Discussion
In this work, we proposed MPIFA, an end-to-end, retraining-free low-rank pruning framework that integrates Pivoting Factorization (PIFA) and an Online Error-Accumulation-Minimization Reconstruction (M) algorithm. PIFA serves as a meta low-rank representation that further compresses existing low-rank decompositions, achieving superior memory savings and inference speedup without additional loss. While this allows PIFA to integrate seamlessly with various low-rank compression techniques, it does not reduce matrix rank on its own and must be applied alongside a low-rank compression method. Meanwhile, M mitigates error accumulation, leading to improved performance. Together, these innovations enable MPIFA to achieve performance comparable to semi-structured pruning while surpassing it in GPU acceleration and compatibility. Future work could explore integrating PIFA into the pretraining stage, as PIFA is fully differentiable, enabling potential efficiency gains during model training.
Acknowledgements
This work was supported by the Zhou Yahui Chair Professorship award of Tsinghua University (to CVC), the National High-Level Talent Program of the Ministry of Science and Technology of China (grant number 20241710001, to CVC).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
(a) FP32 Inference Speedup
(b) FP16 Inference Speedup
(c) FP32 Inference Memory
(d) FP16 Inference Memory
Figure 7: Efficiency of PIFA layer under various ranks, with sequence length = 2048, batch size = 32, and dimension = 8192 on FP32 and FP16 on A6000 GPU. At $50\%$ density, PIFA achieves $47.6\%$ memory savings and 1.95×1.95$\times$1.95 × speedup on FP32. These results guarantee that the overhead of both time and memory of the PIFA layer is quite low.
Figure 8: Condition number. Condition numbers of $\vec{V}^{\mathrm{T}}\mathbf{X}\mathbf{X}^{\mathrm{T}}\vec{V}$ (Equation 5) and $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ (Equation 8) for the first layer of LLaMA2-7B, whose inverses are used in reconstructing $\vec{U}$ and $\vec{V}^{\mathrm{T}}$. Larger calibration sizes reduce condition numbers, enhancing numerical stability and lowering perplexity.
Table 3: Efficiency of PIFA layer and semi-sparse layer across different dimensions, compared to dense linear at same dimension on same GPU, with sequence length of 2048 and batch size of 32, using FP16 (FP32 is not supported by 2:4 sparsity in torch.sparse). The highest speedup and lowest memory are indicated in bold. PIFA shows increasing speedup as the dimension grows. For matrix multiplication with weight matrix shape 32768$\times$32768, cuSPARSELt raises CUDA error.
Dimension
GPU Kernel 32768 16384 8192 4096
Speedup A6000 2:4 (cuSPARSELt) Error 0.94$\times$ 0.97$\times$ 1.09$\times$
2:4 (CUTLASS) 0.79$\times$ 0.92$\times$ 1.15$\times$ 1.18$\times$
PIFA 55% 2.10$\times$ 1.88$\times$ 1.70$\times$ 1.43$\times$
A100 2:4 (cuSPARSELt) Error 1.19$\times$ 1.31$\times$ 1.68$\times$
2:4 (CUTLASS) 1.19$\times$ 1.12$\times$ 1.09$\times$ 1.52$\times$
PIFA 55% 1.98$\times$ 1.70$\times$ 1.54$\times$ 1.37$\times$
Memory 2:4 (cuSPARSELt / CUTLASS) 0.564 0.569 0.589 0.651
PIFA 55% 0.552 0.558 0.578 0.645
Table 4: End-to-end efficiency of MPIFANS on LLaMA2 models, on FP16 (FP32 isn’t supported by semi-sparse). The highest throughput and lowest memory are indicated in bold. MPIFANS consistently outperforms semi-sparse in both throughput and memory with 55% density. Enabling KV cache for semi-sparse model will cause SparseSemiStructuredTensorCUSPARSELT only supports a specific set of operations, can’t perform requested op (expand.default)
Model Metrics GPU Use KV Cache Dense 2:4 (cuSPARSELt) 2:4 (CUTLASS) MPIFANS 55%
llama2-7b Throughput (token/s) A6000 No 354.9 306.6 327.5 472.6
Yes 3409 Error Error 4840
A100 No 614.8 636.2 582.3 822.2
Yes 6918 Error Error 7324
Memory (GB) 12.55 7.274 7.290 7.174
llama2-13b Throughput (token/s) A6000 No 190.0 163.2 180.0 268.7
Yes 2156 Error Error 2721
A100 No 345.4 362.4 321.5 473.3
Yes 4217 Error Error 4532
Memory (GB) 24.36 13.90 13.99 13.69
Appendix A Closed-Form Solution of $\mathbf{V}^{\mathrm{T}}$
We aim to prove that minimizing the Frobenius norm $\|\mathbf{Y}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X}\|_{F}$ with respect to $\mathbf{V}^{\mathrm{T}}$ is equivalent to performing the following two-step optimization:
1. First, minimize $\|\mathbf{Y}-\mathbf{W}\mathbf{X}\|_{F}$ with respect to $\mathbf{W}$.
2. Then, minimize $\|\mathbf{W}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\|_{F}$ with respect to $\mathbf{V}^{\mathrm{T}}$.
We begin by directly minimizing $\|\mathbf{Y}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X}\|_{F}^{2}$ with respect to $\mathbf{V}^{\mathrm{T}}$.
A.1 Direct Optimization
$f(\mathbf{V})$ $=\|\mathbf{Y}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X}\|_{F}^{2}$
$=\operatorname{Tr}\left((\mathbf{Y}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X})^{\mathrm{T}}(\mathbf{Y}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X})\right)$
$=\operatorname{Tr}\left(\mathbf{Y}^{\mathrm{T}}\mathbf{Y}-\mathbf{Y}^{\mathrm{T}}\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X}-\mathbf{X}^{\mathrm{T}}\mathbf{V}\mathbf{U}^{\mathrm{T}}\mathbf{Y}+\mathbf{X}^{\mathrm{T}}\mathbf{V}\mathbf{U}^{\mathrm{T}}\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X}\right)$
$=\operatorname{Tr}(\mathbf{Y}^{\mathrm{T}}\mathbf{Y})-2\operatorname{Tr}(\mathbf{V}^{\mathrm{T}}\mathbf{X}\mathbf{Y}^{\mathrm{T}}\mathbf{U})+\operatorname{Tr}(\mathbf{V}^{\mathrm{T}}\mathbf{X}\mathbf{X}^{\mathrm{T}}\mathbf{V}\mathbf{U}^{\mathrm{T}}\mathbf{U})$
Let us define intermediate matrices:
$\mathbf{A}$ $=\mathbf{X}\mathbf{Y}^{\mathrm{T}}\mathbf{U}$
$\mathbf{B}$ $=\mathbf{X}\mathbf{X}^{\mathrm{T}}$
$\mathbf{C}$ $=\mathbf{U}^{\mathrm{T}}\mathbf{U}$
The objective function becomes:
$f(\mathbf{V})=\text{const}-2\operatorname{Tr}(\mathbf{V}^{\mathrm{T}}\mathbf{A})+\operatorname{Tr}(\mathbf{V}^{\mathrm{T}}\mathbf{B}\mathbf{V}\mathbf{C})$
where ”const” denotes terms independent of $\mathbf{V}$.
Compute the gradient of $f(\mathbf{V})$ with respect to $\mathbf{V}$:
$\frac{\partial f}{\partial\mathbf{V}}=-2\mathbf{A}+2\mathbf{B}\mathbf{V}\mathbf{C}$
Set the gradient to zero:
$-2\mathbf{A}+2\mathbf{B}\mathbf{V}\mathbf{C}=0\implies\mathbf{B}\mathbf{V}\mathbf{C}=\mathbf{A}$
Assuming $\mathbf{B}$ and $\mathbf{C}$ are invertible, we solve for $\mathbf{V}$:
$\mathbf{V}=\mathbf{B}^{-1}\mathbf{A}\mathbf{C}^{-1}$
Substituting back the definitions of $\mathbf{A},\mathbf{B},\mathbf{C}$:
$\mathbf{V}=(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}\left(\mathbf{X}\mathbf{Y}^{\mathrm{T}}\mathbf{U}\right)(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}$
Simplify:
$\mathbf{V}^{\mathrm{T}}=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{Y}\mathbf{X}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}$
A.2 Two-Step Optimization
Now, we perform the two-step optimization and show it leads to the same result.
First, minimize $\|\mathbf{Y}-\mathbf{W}\mathbf{X}\|_{F}^{2}$ with respect to $\mathbf{W}$.
Compute the gradient:
$\frac{\partial}{\partial\mathbf{W}}\|\mathbf{Y}-\mathbf{W}\mathbf{X}\|_{F}^{2}=-2(\mathbf{Y}-\mathbf{W}\mathbf{X})\mathbf{X}^{\mathrm{T}}$
Set the gradient to zero:
$(\mathbf{Y}-\mathbf{W}\mathbf{X})\mathbf{X}^{\mathrm{T}}=0\implies\mathbf{Y}\mathbf{X}^{\mathrm{T}}=\mathbf{W}\mathbf{X}\mathbf{X}^{\mathrm{T}}$
Assuming $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ is invertible:
$\mathbf{W}=\mathbf{Y}\mathbf{X}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}$
Next, minimize $\|\mathbf{W}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\|_{F}^{2}$ with respect to $\mathbf{V}^{\mathrm{T}}$.
Compute the gradient:
$\frac{\partial}{\partial\mathbf{V}}\|\mathbf{W}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\|_{F}^{2}=-2\mathbf{U}^{\mathrm{T}}(\mathbf{W}-\mathbf{U}\mathbf{V}^{\mathrm{T}})$
Set the gradient to zero:
$\mathbf{U}^{\mathrm{T}}\mathbf{W}=\mathbf{U}^{\mathrm{T}}\mathbf{U}\mathbf{V}^{\mathrm{T}}$
Assuming $\mathbf{U}^{\mathrm{T}}\mathbf{U}$ is invertible:
$\mathbf{V}^{\mathrm{T}}=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{W}$
Substitute $\mathbf{W}$:
$\mathbf{V}^{\mathrm{T}}=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}\left(\mathbf{Y}\mathbf{X}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}\right)$
Simplify:
$\mathbf{V}^{\mathrm{T}}=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{Y}\mathbf{X}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}$
A.3 Conclusion
The solution for $\mathbf{V}^{\mathrm{T}}$ obtained through both the direct optimization and the two-step optimization is:
$\mathbf{V}^{\mathrm{T}}=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}\mathbf{Y}\mathbf{X}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}$
Therefore, minimizing $\|\mathbf{Y}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\mathbf{X}\|_{F}$ with respect to $\mathbf{V}^{\mathrm{T}}$ is equivalent to first optimizing $\|\mathbf{Y}-\mathbf{W}\mathbf{X}\|_{F}$ with respect to $\mathbf{W}$, and then optimizing $\|\mathbf{W}-\mathbf{U}\mathbf{V}^{\mathrm{T}}\|_{F}$ with respect to $\mathbf{V}^{\mathrm{T}}$.
Appendix B Experiment Details
B.1 MPIFA
The full method of MPIFA is outlined in Algorithm 3.
Algorithm 3 MPIFA
0:  Original weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$; calibration input from dense $\mathbf{X}_{o}\in\mathbb{R}^{n\times b}$; calibration input from low rank $\mathbf{X}_{u}\in\mathbb{R}^{n\times b}$; target rank $r$; original output ratio $\lambda$
Part 1: Online Error-Accumulation-Minimization Reconstruction
1:  Compute dense output as dense input of next module: $\mathbf{Y}_{o}=\mathbf{W}\mathbf{X}_{o}$
2:  Use SVD-LLM’s pruning (truncation-aware data whitening) to convert to low-rank matrix: $\vec{U},\vec{V}^{\mathrm{T}}\leftarrow\text{SVD-LLM}(\mathbf{W})$
3:  Compute $\vec{X}\vec{X}^{\mathrm{T}}$ accumulatively: $\mathbf{X}\mathbf{X}^{\mathrm{T}}\leftarrow\sum_{i=1}^{b}\mathbf{x}_{u}^{i}{\mathbf{x}_{u}^{i}}^{\mathrm{T}}$
4:  Compute $\vec{Y}_{t}\vec{X}^{\mathrm{T}}$ accumulatively: $\mathbf{Y}_{t}\mathbf{X}^{\mathrm{T}}\leftarrow\sum_{i=1}^{b}(\lambda\mathbf{W}\mathbf{x}_{o}^{i}+(1-\lambda)\mathbf{W}\mathbf{x}_{u}^{i}){\mathbf{x}_{u}^{i}}^{\mathrm{T}}$
5:  Reconstruct $\vec{U}$ as $\mathbf{U}_{r}$: $\mathbf{U}_{r}\leftarrow(\mathbf{Y}_{t}\mathbf{X}^{\mathrm{T}})\vec{V}(\vec{V}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})\vec{V})^{-1}$
6:  Reconstruct $\vec{V}$ as $\mathbf{V}_{r}$: $\mathbf{V}_{r}^{\mathrm{T}}\leftarrow(\mathbf{U}_{r}^{\mathrm{T}}\mathbf{U}_{r})^{-1}\mathbf{U}_{r}^{\mathrm{T}}(\mathbf{Y}_{t}\mathbf{X}^{\mathrm{T}})(\mathbf{X}\mathbf{X}^{\mathrm{T}})^{-1}$
7:  Compute low-rank output as low-rank input of next module: $\mathbf{Y}_{u}=\mathbf{W}\mathbf{X}_{u}$
Part 2: PIFA 8:  Compute low-rank matrix $\mathbf{W}^{\prime}$: $\mathbf{W}^{\prime}\leftarrow\mathbf{U}_{r}\mathbf{V}_{r}^{\mathrm{T}}$ 9:  Use Algorithm 1 to build the PIFA layer $P$ using low-rank matrix $\mathbf{W}^{\prime}$
9:  PIFA layer $P$
A potential issue is that $\mathbf{X}\mathbf{X}^{\mathrm{T}}$ can be singular in some cases, leading to NaN values when calculating the inverse matrix during $\vec{V}$ reconstruction. To address this, we leverage prior knowledge that $\vec{U}\vec{V}^{\mathrm{T}}$ should approximate $\mathbf{W}$ by adding a regularization term to the original optimization target, modifying Equation 8 as follows:
$\mathbf{V}_{r}^{\mathrm{T}}$ $=\arg\min_{\mathbf{V}^{\mathrm{T}}}\|\mathbf{Y}_{t}-\vec{U}\vec{V}^{\mathrm{T}}\mathbf{X}\|_{\text{F}}+\alpha\|\mathbf{W}-\vec{U}\vec{V}^{\mathrm{T}}\|_{\text{F}}$ (9)
$=(\mathbf{U}^{\mathrm{T}}\mathbf{U})^{-1}\mathbf{U}^{\mathrm{T}}(\mathbf{Y}_{t}\mathbf{X}^{\mathrm{T}}+\alpha\mathbf{W})(\mathbf{X}\mathbf{X}^{\mathrm{T}}+\alpha\mathbf{I})^{-1}$
where $\alpha$ is the regularization coefficient, set to 0.001 in all experiments. Regularization is unnecessary for reconstructing $\vec{U}$, as no singularity issues were observed for $\vec{V}^{\mathrm{T}}(\mathbf{X}\mathbf{X}^{\mathrm{T}})\vec{V}$.
B.2 MPIFANS
MPIFANS is the non-uniform sparsity variant of MPIFA, designed to leverage different sparsity distributions across model layers and module types for improved performance. It employs 512 calibration samples. This approach incorporates two key components to define module densities: Type Density and Layer Density, which are combined multiplicatively to determine the final density for each module.
Type Density.
[…]
Layer Density.
Layer Density accounts for non-uniform sparsity across layers. For this, MPIFANS adopts the layerwise density distribution from OWL (Yin et al., ), which identifies layer-wise densities based on outlier distribution. By directly utilizing these precomputed layer densities, MPIFANS ensures that density is allocated more effectively across layers, balancing pruning across regions of varying importance.
Final Module Density.
The final density for each module in MPIFANS is calculated as:
Module Density=Type Density×Layer DensityGlobal $\text{Module Density}=\frac{\text{Type Density}\times\text{Layer Density}}{\text{Global Density}}.$
This formulation ensures that the final density for each module accounts for both type- and layer-specific sparsity requirements, leading to a more effective pruning configuration optimizing performance while maintains the global density same.
In summary, MPIFANS combines the benefits of non-uniform sparsity across both types of modules and individual layers, achieving better performance while ensuring the global density of the model remains unchanged.
B.3 MPIFANS Fine-tuning
Fine-tuning is performed using a mixed dataset comprising the training set of WikiText2 and one shard (1/1024) of the training set of C4 (Raffel et al., 2020). WikiText2’s training set is more aligned with the evaluation dataset but contains a limited number of tokens, whereas the C4 dataset is significantly larger but less aligned with the test set. To balance these characteristics, the datasets are mixed at a ratio of 2% WikiText2 to 98% C4.
We limit fine-tuning to a single epoch, which requires approximately one day on a single GPU (around 1000 steps). The fine-tuning process updates all pruned parameters, including low-rank matrices and semi-sparse matrices, while keeping other parameters, such as embeddings, fixed.
The learning rate is set to $3\times 10^{-6}$, with a warmup phase covering the first 5% of total steps, followed by a linear decay to zero. The sequence length is fixed at 1024, with a batch size of 1 and gradient accumulation steps of 128.
[…]
tableC4 Perplexity ($\downarrow$) at different parameter density (proportion of remaining parameters relative to the original model). The best-performing method is highlighted in bold. Density Model Method 100% 90% 80% 70% 60% 50% 40% LLaMA2-7B SVD 7.29 18931 27154 37208 56751 58451 70567 ASVD 7.98 12.46 201.0 9167 25441 24290 SVD-LLM 13.95 19.89 33.02 61.97 129.8 262.9 MPIFA 8.15 10.20 14.68 25.43 52.01 97.71 LLaMA2-13B SVD 6.74 1994 6301 37250 22783 18196 84680 ASVD 7.15 9.30 23.54 468.5 3537 3703 SVD-LLM 10.93 14.99 24.44 46.65 110.4 267.8 MPIFA 7.27 8.79 21.00 21.33 42.03 80.47 LLaMA2-70B SVD 5.74 10.16 23.28 121.4 1659 7045 12039 ASVD OOM OOM OOM OOM OOM OOM SVD-LLM 7.12 8.71 12.21 21.40 44.10 103.3 MPIFA 6.00 6.76 8.67 13.60 29.04 63.38 LLaMA3-8B SVD 9.47 323597 461991 172968 70896 143573 271176 ASVD 14.43 272.1 8511 18701 108117 9466 SVD-LLM 38.54 98.65 223.5 460.0 784.8 1416 MPIFA 14.76 22.45 44.62 123.0 257.4 429.2
Appendix C Perplexity Evaluation on C4 Dataset
To complement our evaluation on WikiText2, we assess MPIFA on the C4 dataset (Raffel et al., 2020), a widely used benchmark for evaluating LLM perplexity on large-scale, real-world text distributions. For consistency with the main text, we use WikiText2 as the calibration dataset for all methods in this evaluation. The results, summarized in Table B.3, demonstrate that MPIFA consistently achieves the lowest perplexity across most density levels and model sizes, reducing the perplexity gap by 47.6% on LLaMA2-7B, 34.5% on LLaMA2-13B, 55.3% on LLaMA2-70B, and 62.6% on LLaMA3-8B on average across all densities, compared to the best-performing baseline.
Appendix D Zero-Shot Evaluation on SuperGLUE Benchmark
To provide a more comprehensive evaluation of MPIFA’s performance across different compression levels, we conduct zero-shot evaluations at various compression rates on the SuperGLUE benchmark (Wang et al., 2019) using the LLaMA2-7B model.
Evaluations are conducted using the lm-evaluation-harness framework (Gao et al., 2023). All tasks are evaluated using accuracy ($\uparrow$), except for the ReCoRD task, which uses F1 score. The detailed accuracy results are presented in Table 5. The best-performing method at each setting is highlighted in bold. MPIFA consistently achieves the highest mean accuracy across all density levels, outperforming other low-rank methods across a wide range of tasks.
Table 5: Zero-shot evaluations on SuperGLUE datasets at different parameter density on LLaMA2-7B. All tasks are evaluated using accuracy ($\uparrow$), except for the ReCoRD task, which uses F1 score ($\uparrow$). The best-performing method is highlighted in bold.
Density Method BoolQ CB COPA MultiRC ReCoRD RTE WIC WSC Mean
100% Dense 77.7 42.9 87.0 57.0 91.6 63.2 49.7 36.5 63.2
90% SVD 42.6 39.3 67.0 51.0 16.4 55.6 49.8 62.5 48.0
ASVD 55.9 37.5 69.0 47.1 42.5 53.4 49.8 41.3 49.6
SVD-LLM 49.1 41.1 79.0 57.1 87.8 52.7 48.0 48.1 57.8
MPIFA 74.4 64.3 86.0 56.7 91.2 58.5 49.7 36.5 64.7
80% SVD 45.9 57.1 59.0 48.9 12.5 47.3 50.0 58.7 47.4
ASVD 41.6 33.9 58.0 46.8 24.6 55.2 50.0 60.6 46.3
SVD-LLM 44.2 41.1 79.0 55.7 84.7 53.1 50.6 57.7 58.3
MPIFA 69.4 41.1 83.0 55.8 90.3 53.4 48.7 36.5 59.8
70% SVD 40.2 46.4 62.0 43.1 12.4 53.8 48.9 63.5 46.3
ASVD 39.2 46.4 59.0 48.2 14.0 49.1 50.3 63.5 46.2
SVD-LLM 44.4 39.3 82.0 44.6 79.8 53.8 48.9 60.6 56.7
MPIFA 64.8 41.1 80.0 57.2 87.5 53.8 49.1 42.3 59.5
60% SVD 45.5 41.1 61.0 49.4 11.4 51.6 50.0 60.6 46.3
ASVD 48.9 37.5 59.0 46.6 15.2 50.2 50.0 40.4 43.5
SVD-LLM 38.5 39.3 71.0 44.9 66.4 53.4 50.2 59.6 52.9
MPIFA 56.6 35.7 79.0 44.5 82.8 57.8 52.0 63.5 59.0
50% SVD 38.6 44.6 63.0 43.0 10.1 52.7 50.2 63.5 45.7
ASVD 42.2 39.3 63.0 45.7 14.4 52.7 50.0 63.5 46.3
SVD-LLM 37.9 41.1 66.0 42.8 50.5 52.7 50.0 63.5 50.5
MPIFA 38.0 35.7 70.0 42.8 71.1 52.4 50.0 63.5 52.9
40% SVD 49.5 42.9 61.0 48.1 11.3 52.0 50.0 49.0 45.5
ASVD 42.6 42.9 59.0 47.7 14.5 53.1 50.0 64.4 46.8
SVD-LLM 37.8 41.1 67.0 42.8 37.0 52.7 50.0 63.5 49.0
MPIFA 37.8 37.5 68.0 42.8 54.7 53.8 50.0 63.5 51.0
Appendix E Comparison with Structured Pruning Baseline
To provide a fair comparison with structured pruning methods, we include additional evaluations of LLM-Pruner (Ma et al., 2023) as a baseline. All experiments are conducted on the WikiText2 dataset using the LLaMA2-7B model.
Perplexity Comparison.
Table 6 shows the perplexity across various parameter densities. On average, MPIFA reduces the perplexity gap by 87.2% compared to LLM-Pruner.
Table 6: LLM-Pruner vs. MPIFA: Perplexity Comparison on WikiText2 using LLaMA2-7B at different densities. MPIFA consistently achieves lower perplexity across all densities.
Method 100% 90% 80% 70% 60% 50% 40%
LLM-Pruner 5.47 6.58 8.81 13.70 40.49 126.0 1042
MPIFA 5.69 6.16 7.05 8.81 12.77 21.25
Inference Speedup and Memory Efficiency.
Table 7 and Table 8 compare the inference speedup and memory usage (relative to dense linear layer) for PIFA and LLM-Pruner layers, across different hidden dimensions on an A6000 GPU.
Table 7: Inference speedup (× over dense) for PIFA and LLM-Pruner layers.
Method (density) d=16384 d=8192 d=4096
PIFA (55%) 1.88× 1.70× 1.43×
LLM-Pruner (55%) 1.81× 1.77× 1.67×
LLM-Pruner (70%) 1.42× 1.41× 1.35×
Table 8: Memory usage (× over dense) for PIFA and LLM-Pruner layers.
Method (density) d=16384 d=8192 d=4096
PIFA (55%) 0.56× 0.58× 0.64×
LLM-Pruner (55%) 0.56× 0.58× 0.65×
LLM-Pruner (70%) 0.70× 0.72× 0.75×
At the same density (55%), PIFA achieves similar speedup and memory efficiency as LLM-Pruner. When comparing MPIFA at 55% density to LLM-Pruner at 70% density, MPIFA consistently achieves lower perplexity, faster inference, and reduced memory usage.
Appendix F Compression Time and Memory Usage Comparison
We provide a detailed comparison of the compression time and peak GPU memory usage during compression across different methods. These metrics reflect the practical resource requirements of each approach.
Compression Time.
Table 9 reports the time required to perform compression for each method, measured on a single A6000 GPU. On an A100 GPU, the compression time is approximately halved. ASVD requires significantly longer compression time, up to 10–20 hours, due to the need for sensitivity-based truncation rank searching. The M method involves only reconstruction.
Table 9: Compression time for different methods on a single A6000 GPU.
Model ASVD SVD-LLM PIFA M
LLaMA2-7B 10h 30 min 15 min 30 min
LLaMA2-13B 20h 1h 30 min 1h
Peak Memory Usage During Compression.
Table 10 reports the peak GPU memory usage during compression. Method M has a relatively low memory footprint, as it processes one layer and one sample at a time.
Table 10: Peak GPU memory usage (GB) during compression for different methods.
Model ASVD SVD-LLM PIFA M
LLaMA2-7B 15G 20G 0.5G 6G
LLaMA2-13B 30G 25G 1G 10G
For the M method, we report only the reconstruction time, excluding the time required for the initial low-rank pruning step, which could be SVD or SVD-LLM. The low memory footprint of M is primarily attributed to:
  • Online calibration: Only the current sample is loaded into GPU memory, while the rest of the samples remain in CPU memory until needed.
  • Layer-wise loading: Only the current pruning layer is loaded to the GPU at any time, with all other layers remaining on CPU.
Appendix G Enhancing Other Low-Rank Pruning Methods with PIFA and M
To demonstrate the generality of our proposed methods, we evaluate how PIFA and M can be applied on top of existing low-rank pruning techniques beyond SVD-LLM, including the pruning strategies proposed in the ESPACE paper (Sakr & Khailany, ).
For a pruning-only comparison, we reproduce the pruning step of ESPACE and compare its performance when combined with PIFA and M. ESPACE introduces six variants: MSE (Eq. 6), MSE-NORM (Eq. 7), GO-MSE (Eq. 8), GO-MSE-NORM (Eq. 8), NL-MSE (Eq. 9), and NL-MSE-NORM (Eq. 9). We exclude the NL-MSE variants as they rely on backpropagation, which is infeasible on memory-constrained GPUs.
Table 11 reports the perplexity (PPL) on WikiText2 at 50% density using the LLaMA2-7B model, showing how PIFA and M further improve various low-rank pruning baselines.
Table 11: Perplexity (PPL) on WikiText2 at 50% density using LLaMA2-7B.
Pruning Method (X) X X + PIFA X + M X + MPIFA
SVD-LLM (W) 33.27 19.64 16.55 12.77
ESPACE (MSE) 280.19 144.32 20.99 16.84
ESPACE (MSE-NORM) 172.30 113.82 21.73 17.42
ESPACE (GO-MSE) 41.75 24.17 17.47 13.55
ESPACE (GO-MSE-NORM) 37.45 23.19 17.47 13.63
Here, SVD-LLM (W) refers to the standalone pruning output from SVD-LLM without additional reconstruction.
PIFA and M consistently improve the performance of different low-rank pruning methods, including ESPACE, demonstrating that they are general-purpose techniques applicable across various low-rank strategies. These results suggest that PIFA and M are not specific to SVD-LLM but can serve as plug-in modules for other pruning methods to improve accuracy while maintaining efficiency.